# Finely Stratified Rerandomization Designs

### v2

@@ 2 Framework and Designs
In the exactly identified case, $\widehat{\theta}$ solves $\widehat{g}(\widehat{\theta})=0$. In the next section, we study generalized method of moments (GMM) estimation of the causal parameters $\theta_{0}$ and $\theta_{n}$ under stratified rerandomization.
Remark 2.9 (M-estimation).
Our results below also extend to M-estimators of the form $\widehat{\theta}=\operatorname*{argmax}_{\theta}E_{n}[m(D_{i},R_{i},S_{i},\theta)]$, even when $\widehat{\theta}$ cannot be formulated as a GMM estimator e.g. due to the existence of local maxima. For example, this happens in some nonconvex problems in density estimation (Newey and McFadden, (1994)). We briefly discuss this extension in Section 8.4.
@@ 3 Asymptotics for GMM Estimation
In this section, we characterize the asymptotic distribution of the GMM estimator $\widehat{\theta}$ under stratified rerandomization designs, as in Definition 2.1. We show that the variance under stratified rerandomization is proportional to the residuals of a partially linear regression model, up to an extra term that reflects slackness in the rerandomization criterion. In this sense, stratified rerandomization does partially linear regression adjustment “by design.” First, we state some technical conditions that are needed for the following results.
@@
Orthogonal Expansion. Recall $\phi(W,\theta)=E[g(D,R,S,\theta)|W]$ for $W=(R,S(1),S(0))$. Define the assignment influence component $a(W,\theta)\equiv\operatorname{Var}(D)(g_{1}(W,\theta)-g_{0}(W,\theta))$. For Horvitz-Thompson weights $H=(D-p)/(p-p^{2})$, a simple calculation shows that we can expand
@@
Our work below shows that $a(W,\theta)$ parameterizes estimator variance due to random assignment, while $\phi(W,\theta)$ parameterizes the variance due to random sampling. We work directly with this expansion in what follows.
Example 3.3 (SATE).
Continuing Example 2.6 above, let $\bar{Y}=(1-p)Y(1)+pY(0)$, a convex combination that summarizes each unit’s potential outcome level. Then for the score $g(D,Y,\theta)=HY-\theta$, we have $a(W,\theta)=\bar{Y}$. A simple calculation shows that for $\widehat{\theta}=E_{n}[H_{i}Y_{i}]$ and $\theta_{n}=E_{n}[Y_{i}(1)-Y_{i}(0)]$, we have
$\widehat{\theta}-\theta_{n}=E_{n}[H_{i}a(W_{i})]=\frac{\operatorname{Cov}_{n}(D_{i},\bar{Y}_{i})}{\operatorname{Var}_{n}(D_{i})}.$ (3.2)
Intuitively, the term $E_{n}[H_{i}a(W_{i})]$ from Equation 3.1 isolates the estimator variance due to chance in-sample correlations between the assignments $D_{i}$ and outcome levels $\bar{Y}_{i}$. By contrast, $\phi(W,\theta)=Y(1)-Y(0)-\theta$ does not depend on assignments $D_{i}$, and $\operatorname{Var}(\phi(W,\theta))=\operatorname{Var}(Y(1)-Y(0))$ isolates the estimator variance due to random sampling of units and heterogeneity of individual treatment effects.
@@ 3.1 Finite Population Estimand
Our first theorem studies GMM estimation of the finite population estimand $\theta_{n}$, which solves $E_{n}[\phi(W_{i},\theta_{n})]=0$. We extend these results to $\theta_{0}$ in Corollary 3.8 below. To state the theorem, define the GMM linearization matrix $\Pi=-(G^{\prime}MG)^{-1}G^{\prime}M\in\mathbb{R}^{d_{\theta}\times d_{g}}$. Note that in the exactly identified case $d_{g}=d_{\theta}$, we just have $\Pi=-G^{-1}$. For brevity, we also denote $v_{D}=\operatorname{Var}(D)=p-p^{2}$.
@@
Lemma 3.4 generalizes Equation 3.2 above, showing that
$\widehat{\theta}-\theta_{n}=\frac{\operatorname{Cov}_{n}(D_{i},\Pi a(W_{i},\theta_{0}))}{\operatorname{Var}_{n}(D_{i})}+o_{p}(n^{-1/2}).$
This implies that to first order, the errors in estimating $\theta_{n}$ are driven by the random in-sample correlations between treatment assignments $D_{i}$ and the assignment influence function $\Pi a(W_{i},\theta_{0})$. Our main theorem shows that, by balancing $\psi$ and $h$, stratified rerandomization reduces these correlations, improving precision.
@@
Let $\gamma_{0}$ be optimal in Equation 3.3. The term $R_{A}$ is a truncated Gaussian
@@
Note that the variance matrix $V_{a}\in\mathbb{R}^{d_{\theta}\times d_{\theta}}$, so the minimum should be interpreted in the positive semidefinite sense. In particular, we say $V(\gamma_{0})=\min_{\gamma}V(\gamma)$ if $V(\gamma_{0})\preceq V(\gamma)$ for all $\gamma\in\mathbb{R}^{d_{h}\times d_{\theta}}$. Theorem 3.5 shows that $\sqrt{n}(\widehat{\theta}-\theta_{n})$ is asymptotically distributed as an independent sum of a normal $\mathcal{N}(0,V_{a})$ and truncated normal $R_{A}$. The normal term $\mathcal{N}(0,V_{a})$ only depends on the “treatment assignment” component of the influence function, $\Pi a(W,\theta_{0})$. The variance is attenuated nonparametrically by the stratification variables $\psi$ and linearly by rerandomization covariates $h$.
@@
To isolate the precision gains due to rerandomization, the following corollary specializes Theorem 3.5 to the case of stratification without rerandomization ($A=\mathbb{R}^{d_{h}}$), as well as complete randomization, as defined in Examples 2.3 and 2.4.
@@
Suppose $D_{1:n}$ as in Definition 2.1 with $A=\mathbb{R}^{d_{h}}$. Require Assumption 3.1. Then $\sqrt{n}(\widehat{\theta}-\theta_{n})|W_{1:n}\Rightarrow\mathcal{N}(0,V)$ with $V=v_{D}^{-1}E[\operatorname{Var}(\Pi a(W,\theta_{0})|\psi)]$. In particular, if $D_{1:n}\sim\operatorname{CR}(p)$ then $V=v_{D}^{-1}\operatorname{Var}(\Pi a(W,\theta_{0}))$.
Corollary 3.6 shows that fine stratification reduces the variance of GMM estimation to $V=v_{D}^{-1}E[\operatorname{Var}(\Pi a(W,\theta_{0})|\psi)]\leq v_{D}^{-1}\operatorname{Var}(\Pi a(W,\theta_{0}))$, a nonparametric improvement. Rerandomization as in Definition 2.1 provides a further linear variance reduction to $V_{a}=\min_{\gamma\in\mathbb{R}^{d_{h}\times d_{\theta}}}E[\operatorname{Var}(\Pi a(W,\theta_{0})-\gamma^{\prime}h|\psi)]$, up to the residual imbalance term $R_{A}$.
Remark 3.7 (Design-Based Asymptotics).
Our results above show that $\sqrt{n}(\widehat{\theta}-\theta_{n})|W_{1:n}\Rightarrow\mathcal{N}(0,V_{a})+R_{A}$, conditional on the sampled data $W_{1:n}=(R_{i},S_{i}(1),S_{i}(0))_{i=1}^{n}$. [^15] See Proposition 8.16 in the appendix for a formal statement. This result is “design-based” in the sense that the variance in the limiting distribution arises solely due to randomness of the treatment assignments $D_{1:n}$. However, we impose structure on the sequence of populations $W_{1:n}$ ex-ante, assuming each population is drawn from a fixed measure, $W_{i}\sim F$. This allows us to provide simple variance expressions that connect our results with the superpopulation-based literature on GMM and partially linear adjustment in econometrics. By contrast, the “sequence of finite populations model” often used in the statistics literature (e.g. Li et al., (2018)) begins with an arbitrary sequence of finite populations $(W_{i,n})_{i=1}^{n}$, imposing the minimal structure needed for certain moments to converge ex-post. It would be interesting to extend our results to this setting, but we leave this to future work.
@@ 3.2 Superpopulation Estimand
The next result extends Theorem 3.5 to the superpopulation estimand $\theta_{0}$, which uniquely solves $E[\phi(W,\theta_{0})]=0$.
@@
Comparing Corollary 3.8 with the results above, we see that targeting $\theta_{0}$ instead of $\theta_{n}$ adds an extra independent Gaussian term $\mathcal{N}(0,V_{\phi})$ to the asymptotic distribution. Intuitively, $V_{\phi}$ arises due to iid random sampling of $\Pi\phi(W,\theta_{0})$. Notice that stratification and rerandomization only affect the assignment influence function component $\Pi a(W,\theta_{0})$, while the sampling influence component $\Pi\phi(W,\theta_{0})$ is irreducible. In this sense, the statistical consequences of different designs and adjustment strategies all happen at the level of the finite population estimand $\theta_{n}$, while targeting the superpopulation estimand $\theta_{0}$ just adds extra irreducible noise. For pure stratification, Bai et al., (2024) were the first to derive an analogue of part (b) of Corollary 3.8 in the exactly identified case, under different GMM regularity conditions than we use here.
@@
Continuing Example 2.6, we had $\phi(W,\theta)=Y(1)-Y(0)-\theta$, so $G=1$ and $\Pi=1$. As above, $a(W,\theta)=(1-p)Y(1)+pY(0)\equiv\bar{Y}$. The GMM estimator $\widehat{\theta}=\bar{Y}_{1}-\bar{Y}_{0}$ is just difference of means. Then by Theorem 3.5 and Corollary 3.8, we have $\sqrt{n}(\widehat{\theta}-\operatorname{SATE})|W_{1:n}\Rightarrow\mathcal{N}(0,V_{a})+R_{A}$ and $\sqrt{n}(\widehat{\theta}-\operatorname{ATE})\Rightarrow\mathcal{N}(0,V_{\phi}+V_{a})+R_{A}$ with
@@
The term $V_{\phi}$, which only appears when estimating the superpopulation estimand $\theta_{0}$, reflects sampling variance due to treatment effect heterogeneity. The term $V_{a}$ is the variance due to random assignment, caused by random in-sample correlations between treatments $D$ and outcome levels $\bar{Y}$. Covariate-adaptive randomization and adjustment can be used to reduce $V_{a}$, while $V_{\phi}$ is an irreducible sampling variance.
@@
Wang et al., (2021) study SATE estimation under stratified rerandomization in the sequence of finite populations framework. Relative to Wang et al., (2021), by imposing the tight-matching condition 2.1 we are able to derive a simple closed form for the asymptotic variance in terms of the measure $W\sim F$, showing an equivalence with partially linear regression adjustment.
Example 3.11 (Treatment Effect Heterogeneity).
Continuing Example 2.7, consider the case with perfect compliance $D=Z$ and $f(X,\theta)=X^{\prime}\theta$. Then we can use the slightly modified score $g(D,X,Y,\theta)=(HY-X^{\prime}\theta)X$. Then for $\tau=Y(1)-Y(0)$ we have $\phi(W,\theta_{0})=(\tau-X^{\prime}\theta_{0})X$, and the parameters $\theta_{n}$, $\theta_{0}$ are the best linear predictors of treatment effect heterogeneity
@@
It’s also easy to see that $a(W,\theta_{0})=\bar{Y}X$ and $\Pi=E[XX^{\prime}]^{-1}$. Then for $e=\tau-X^{\prime}\theta_{0}$, the variance matrices in Corollary 3.8 are
$V_{\phi}=E[XX^{\prime}]^{-1}E[e^{2}XX^{\prime}]E[XX^{\prime}]^{-1}\quad\quad V_{a}=\min_{\gamma\in\mathbb{R}^{d_{h}\times d_{x}}}v_{D}^{-1}E[\operatorname{Var}(\bar{Y}\Pi X-\gamma^{\prime}h|\psi)].$
The expression for $V_{a}$ shows that if we want to precisely estimate treatment effect heterogeneity, it is important to stratify and rerandomize not only the variables that predict outcome levels $\bar{Y}$, but also their interactions with the heterogeneity variable $X$.
@@ 3.3 Equivalence with Partially Linear Adjustment
More generally, Theorem 3.5 shows that under stratified rerandomization designs, the usual GMM estimator $\widehat{\theta}$ behaves like semiparametrically adjusted GMM in the iid setting. Formally, let $\mathcal{L}(\psi)=L_{2}^{d_{\theta}}(\psi)$ be the $d_{\theta}$-fold Cartesian product of $L_{2}(\psi)$, the space of square-integrable functions. Then the variance due to random assignment $V_{a}$ in Theorem 3.5 is can be written in terms of the residuals of the influence function $\Pi a(W,\theta_{0})$ in a partially linear regression on $\psi$ and $h$:
@@
For example, for the $\operatorname{SATE}$ estimation problem one can show that $\widehat{\theta}^{*}$ is just an oracle version of the usual augmented inverse propensity weighting (AIPW) estimator (Robins and Rotnitzky, (1995)), with partially linear regression models in each arm. [^16] Feasible partially linear adjustment in an iid mean estimation problem with missing data was studied in Wang et al., (2004). See also the related semiparametric adjustment for GMM parameters in Graham, (2011).
@@
Suppose that $D_{1:n}\sim\operatorname{CR}(p)$. The oracle partially linearly adjusted GMM estimator $\sqrt{n}(\widehat{\theta}^{*}-\theta_{n})|W_{1:n}\Rightarrow\mathcal{N}(0,V_{a})$, with variance $V_{a}$ as defined in Theorem 3.5.
Under a completely randomized design, we require ex-post semiparametric adjustment to achieve $V_{a}$. Under stratified rerandomization, however, the simple GMM estimator $\widehat{\theta}$ automatically achieves $V_{a}$, up to the leftover imbalance term $R_{A}$.

### v3

@@ 2 Framework and Designs
We are mostly interested in the exactly identified case, where $\widehat{\theta}$ solves $\widehat{g}(\widehat{\theta})=0$. In what follows, we study the properties of generalized method of moments (GMM) estimation of the causal parameters $\theta_{0}$ and $\theta_{n}$ under stratified rerandomization.
@@ 3 Asymptotics for GMM Estimation
In this section, we characterize the asymptotic distribution of the GMM estimator $\widehat{\theta}$ under the stratified rerandomization designs in Definition 2.1. We show that the asymptotic variance of $\widehat{\theta}$ is proportional to the residuals of a partially linear regression model, up to a remainder term due to slackness in the rerandomization criterion. In this sense, stratified rerandomization does partially linear regression adjustment “by design.” First, we state some technical conditions that are needed for the following results.
@@
Sampling and Assignment Expansion. Recall $\phi(W,\theta)=E[g(D,R,S,\theta)|W]$ for $W=(R,S(1),S(0))$. Define $a(W,\theta)\equiv\operatorname{Var}(D)(g_{1}(W,\theta)-g_{0}(W,\theta))$, which we refer to as the “assignment function.” For Horvitz-Thompson weights $H=(D-p)/(p-p^{2})$, a calculation shows we can expand
@@
Our results below show that $a(W,\theta)$ parameterizes estimator variance due to random assignment, while $\phi(W,\theta)$ parameterizes the variance due to random sampling for the superpopulation estimand $\theta_{0}$. We work directly with this expansion in what follows.
Example 3.3 (ATE and SATE).
Continuing Example 2.7 above, define $\bar{Y}=(1-p)Y(1)+pY(0)$. This is a convex combination that summarizes both potential outcomes, which we view as the unit’s “outcome level.” Then for the score $g(D,Y,\theta)=HY-\theta$, we have $a(W,\theta)=\operatorname{Var}(D)(Y(1)/p-(-Y(0)/(1-p)))=\bar{Y}$. Another simple calculation [^16] Note that for stratified designs $E_{n}[D_{i}]=p$, so $E_{n}[H_{i}Y_{i}]=\bar{Y}_{1}-\bar{Y}_{0}$. This is not true for iid designs. shows that for difference of means $\widehat{\theta}=E_{n}[H_{i}Y_{i}]$ and estimands $\theta_{n}=\operatorname{SATE}$, $\theta_{0}=\operatorname{ATE}$
$\widehat{\theta}-\theta_{0}$ $=(\widehat{\theta}-\theta_{n})+(\theta_{n}-\theta_{0})=E_{n}[H_{i}a(W_{i})]+E_{n}[\phi(W_{i},\theta_{0})]$
$=(E_{n}[\bar{Y}_{i}|D_{i}=1]-E_{n}[\bar{Y}_{i}|D_{i}=0])+(E_{n}[Y_{i}(1)-Y_{i}(0)]-\theta_{0})$
The assignment term $E_{n}[H_{i}a(W_{i})]$ from Equation 3.1 isolates the estimation error due to chance imbalances in the outcome levels $\bar{Y}_{i}$ between treatment and control during random assignment. By contrast, the term $E_{n}[\phi(W_{i},\theta_{0})]$ for sampling function $\phi(W,\theta)=Y(1)-Y(0)-\theta$ isolates the estimation error due to random sampling of heterogeneous units. The next section shows that these two sources of error are orthogonal. Note also that the error $\widehat{\theta}-\theta_{n}$ for estimating $\theta_{n}$ is only due to assignment imbalances, not sampling variability.
@@ 3.1 Finite Population Estimand
Our first theorem studies GMM estimation of the finite population estimand $\theta_{n}$, which solves $E_{n}[\phi(W_{i},\theta_{n})]=0$. We extend these results to $\theta_{0}$ in Corollary 3.7 below. To state the theorem, define the GMM linearization matrix $\Pi=-(G^{\prime}MG)^{-1}G^{\prime}M\in\mathbb{R}^{d_{\theta}\times d_{g}}$. Note that in the exactly identified case $d_{g}=d_{\theta}$, we just have $\Pi=-G^{-1}$. For brevity, we also denote the recurring constant $v_{D}=\operatorname{Var}(D)=p-p^{2}$.
@@
Lemma 3.4 generalizes Example 3.3 above, showing that
$\widehat{\theta}-\theta_{n}=\Pi\left(E_{n}[a(W_{i},\theta_{0})|D_{i}=1]-E_{n}[a(W_{i},\theta_{0})|D_{i}=0]\right)+o_{p}(n^{-1/2}).$
This implies that the errors in estimating any finite population GMM parameter $\theta_{n}$ are driven by random imbalances in the assignment function $a(W_{i},\theta_{0})$ between treatment and control units, at least to first-order. Our main theorem shows that, by balancing $\psi$ and $h$ ex-ante, stratified rerandomization reduces these imbalances, improving precision.
@@
Let $\gamma_{0}$ be optimal in Equation 3.2. The term $R_{A}$ is a truncated Gaussian vector
@@
Note the variance $V_{a}$ is a matrix $V_{a}\in\mathbb{R}^{d_{\theta}\times d_{\theta}}$, so the minimum should be interpreted in the positive semidefinite sense. In particular, we say $V(\gamma_{0})=\min_{\gamma}V(\gamma)$ if $V(\gamma_{0})\preceq V(\gamma)$ for all $\gamma\in\mathbb{R}^{d_{h}\times d_{\theta}}$. Theorem 3.5 shows that $\sqrt{n}(\widehat{\theta}-\theta_{n})$ is asymptotically distributed as an independent sum of a normal $\mathcal{N}(0,V_{a})$ and truncated normal vector $R_{A}$. Both terms only depend on the “assignment” component of the influence function, $\Pi a(W,\theta_{0})$. The variance is attenuated nonparametrically by the stratification variables $\psi$ and linearly by the rerandomization covariates $h$.
@@
To isolate the precision gains due to rerandomization, the following corollary specializes Theorem 3.5 to the case of stratification without rerandomization ($A=\mathbb{R}^{d_{h}}$), as well as complete randomization, defined in Examples 2.2 and 2.3.
@@
Suppose $D_{1:n}$ as in Definition 2.1 with $A=\mathbb{R}^{d_{h}}$. Require Assumption 3.1. Then $\sqrt{n}(\widehat{\theta}-\theta_{n})|W_{1:n}\Rightarrow\mathcal{N}(0,V_{a})$ with $V_{a}=v_{D}^{-1}E[\operatorname{Var}(\Pi a(W,\theta_{0})|\psi)]$. In particular, if $D_{1:n}$ is completely randomized $\psi=1$, then $V_{a}=v_{D}^{-1}\operatorname{Var}(\Pi a(W,\theta_{0}))$.
Corollary 3.6 shows that fine stratification reduces the variance of GMM estimation of $\theta_{n}$ to $V_{a}=v_{D}^{-1}E[\operatorname{Var}(\Pi a(W,\theta_{0})|\psi)]\leq v_{D}^{-1}\operatorname{Var}(\Pi a(W,\theta_{0}))$, a nonparametric improvement. Rerandomization as in Definition 2.1 provides a further linear variance reduction to $V_{a}=\min_{\gamma\in\mathbb{R}^{d_{h}\times d_{\theta}}}E[\operatorname{Var}(\Pi a(W,\theta_{0})-\gamma^{\prime}h|\psi)]$, up to the residual imbalance term $R_{A}$.
@@ 3.2 Superpopulation Estimand
This section extends the asymptotics above to the superpopulation estimand $\theta_{0}$ solving $E[\phi(W,\theta_{0})]=0$. We show that by targeting $\theta_{0}$ we incur additional sampling variance that is invariant to the distribution of treatment assignments $D_{1:n}$.
@@
The corollary shows that targeting $\theta_{0}$ instead of $\theta_{n}$ adds an extra independent $\mathcal{N}(0,V_{\phi})$ term to the asymptotic distribution. The variance $V_{\phi}$ arises due to iid random sampling of the sampling function $\Pi\phi(W,\theta_{0})$. Notice that stratified rerandomization only reduces the variance due to imbalances in the assignment function $\Pi a(W,\theta_{0})$, while the variance due to sampling $\Pi\phi(W,\theta_{0})$ is irreducible. In this sense, the statistical consequences of different designs and adjustment strategies all happen at the level of the finite population estimand $\theta_{n}$, while targeting the superpopulation $\theta_{0}$ just adds extra sampling noise. Note that for pure stratification, Bai et al., 2024b were the first to derive an analogue of part (b) of Corollary 3.7, under different GMM regularity conditions than we use here. [^18] In particular, Bai et al., 2024b allow for non-smooth GMM scores and impose a VC dimension condition on $g_{d}(W,\theta)$. We restrict to the smooth case, using compactness of $\Theta$ to avoid entropy conditions.
@@
Continuing Example 3.3, Theorem 3.5 and Corollary 3.7 show that $\sqrt{n}(\widehat{\theta}-\operatorname{SATE})|W_{1:n}\Rightarrow\mathcal{N}(0,V_{a})+R_{A}$ and $\sqrt{n}(\widehat{\theta}-\operatorname{ATE})\Rightarrow\mathcal{N}(0,V_{\phi}+V_{a})+R_{A}$ with
@@
The term $V_{\phi}$ reflects sampling variance due to treatment effect heterogeneity. The term $V_{a}$ is the variance due to random assignment, caused by random imbalances in outcome levels $\bar{Y}$ between $D_{i}=1$ and $D_{i}=0$. Covariate-adaptive randomization and adjustment can be used to reduce $V_{a}$, while $V_{\phi}$ is an irreducible sampling variance.
@@
Wang et al., (2021) study SATE estimation under stratified rerandomization in the sequence of finite populations framework. Relative to that work, here we allow for data-adaptive strata $s=s(\psi_{1:n},\pi_{n})$, endogenizing the process of fine stratification. By imposing the tight-matching condition 2.1, which can be satisfied by the matching algorithms in Bai et al., (2021) and Cytrynbaum, 2024b , we are able to derive a simple closed form for the asymptotic variance, providing a novel connection between stratified rerandomization and partially linear regression adjustment.
Example 3.10 (CATE).
Specializing Example 2.8, consider estimating the best linear predictor of treatment effect heterogeneity in an experiment with perfect compliance. We can use the slightly simpler score $g(D,X,Y,\theta)=(HY-X^{\prime}\theta)X$. Then for $\tau=Y(1)-Y(0)$ we have $\phi(W,\theta_{0})=(\tau-X^{\prime}\theta_{0})X$, and the parameters $\theta_{n}$ and $\theta_{0}$ are
@@
The parameter $\theta_{n}$ was studied in Ding et al., (2019) under complete randomization. A simple calculation shows that assignment function $a(W,\theta_{0})=\bar{Y}X$ and $\Pi=E[XX^{\prime}]^{-1}$. Then for residual $e=\tau-X^{\prime}\theta_{0}$, the variances in Corollary 3.7 are
$V_{\phi}=\operatorname{Var}(\Pi eX),\quad\quad V_{a}=\min_{\gamma\in\mathbb{R}^{d_{h}\times d_{x}}}v_{D}^{-1}E[\operatorname{Var}(\Pi\bar{Y}X-\gamma^{\prime}h|\psi)].$
The expression for $V_{a}$ shows that if we want to precisely estimate $\theta_{n}$ and $\theta_{0}$, it is important to include not only the variables that predict outcome levels $\bar{Y}$ in $\psi$ and $h$, but also their interactions with the desired heterogeneity variable $X$. We consider such interacted designs for estimating treatment effect heterogeneity in our simulations and empirical application to Angrist et al., (2013) below.
@@ 3.3 Equivalence with Partially Linear Adjustment
More generally, Theorem 3.5 shows that under stratified rerandomization designs, the unadjusted GMM estimator $\widehat{\theta}$ automatically behaves like semiparametrically adjusted GMM in the completely randomized setting. Formally, let $\mathcal{L}(\psi)=L_{2}^{d_{\theta}}(\psi)$ be the $d_{\theta}$-fold Cartesian product of $L_{2}(\psi)$, the space of square-integrable functions. Then the variance due to random assignment $V_{a}$ in Theorem 3.5 is can be written in terms of the residuals of the influence function $\Pi a(W,\theta_{0})$ in a partially linear regression on $\psi$ and $h$:
@@
For the $\operatorname{SATE}$ estimation problem one can show that $\widehat{\theta}^{*}$ is just an oracle version of the usual augmented inverse propensity weighting (AIPW) estimator (Robins and Rotnitzky, (1995)), with partially linear regression models in each arm. [^19] Feasible partially linear adjustment in an iid mean estimation problem with missing data was studied in Wang et al., (2004). See also the related semiparametric adjustment for GMM parameters in Graham, (2011).
@@
Suppose that $D_{1:n}$ is completely randomized. The oracle partially linearly adjusted GMM estimator $\sqrt{n}(\widehat{\theta}^{*}-\theta_{n})|W_{1:n}\Rightarrow\mathcal{N}(0,V_{a})$, with variance $V_{a}$ as defined in Theorem 3.5.
Under a completely randomized design, we require ex-post semiparametric adjustment to achieve $V_{a}$. Under stratified rerandomization, however, the simple GMM estimator $\widehat{\theta}$ automatically achieves $V_{a}$, up to the residual imbalance term $R_{A}$.